OpenStudy (anonymous):

Help :( I have been stuck on this problem and can't seem to get it: A photon with wavelength λ = 0.1440nm is incident on an electron that is initially at rest. If the photon scatters at an angle of 60.0 ∘ from its original direction, what is the magnitude of the linear momentum of the electron just after the collision with the photon?What is the direction of the linear momentum of the electron just after the collision with the photon? I am using the p=h/lambda and the comptant scatting equation and not getting the right results. Is it correct that angle=tan-1(p photon after/p photon before)

OpenStudy (aaronq):

you used this? http://hyperphysics.phy-astr.gsu.edu/hbase/quantum/compeq.html#c1

OpenStudy (anonymous):

You have to use Compton scattering and momentum \[\lambda ' - \lambda = \lambda_c(c - \cos \phi) \] and \[p = \frac{ h }{ \lambda }\]

OpenStudy (anonymous):

\[\lambda ' = \lambda + \left( \frac{ h }{ mc } \right)(1-\cos \phi)\] then from here you can use conservation of momentum to find the angle

OpenStudy (anonymous):

You need it for both x and y components

OpenStudy (anonymous):

Yeah, I am looking for the angle of the electron and it's momentum though. Is it correct that the tan inverse of the final potential energy of the photon over its initial energy gives the angle of the electron? \[P'/P=\tan \phi \]

OpenStudy (anonymous):

If that's how you've labelled your components, its just like how you did momentum problems back in the day, you use Pythagorus for the momentum of the electron and tan theta for for the direction.

OpenStudy (anonymous):

So if you can show me your set up, it would be easier for me to see what exactly you're doing

OpenStudy (anonymous):

\[P=(hc/\lambda) (1/C) = 1240eV/.144nm = 8611eV/C \] \[\lambda' -.144 = (1240/(.511*10^6))(1-\cos60) ........\lambda' = 0.145nm\] \[P'=1240/0.145=8539eV/c\] \[P'/P=(Pelectronsin \phi)/(Pelectron \cos \phi)=\tan\] \[\phi=44.78\]degrees but that isn't right :( Also thank you for the help I have been stuck on this for a long time

OpenStudy (anonymous):

Nice try! Alright, so lets go over this problem.

OpenStudy (anonymous):

So we have the equations of Compton scattering and of the momentum. So what we have to do here is find the scattered wavelength first, as I showed above, \[\lambda ' = \lambda + \left( \frac{ h }{ mc } \right) (1-\cos \phi)\] so we need this scattered wavelength before we can use conservation of momentum. But, remember as I was saying before we need to break it into components if we want to use conservation of momentum. So we have \[\huge P_{photon}+ P_{electron} = P'_{photon}+P'_{electron}\] momentum conservation. So we'll have it for the x direction and the y direction, can you try that out?

OpenStudy (anonymous):

\[\lambda'=\lambda+hc/mc^2 (1-\cos60) \] I multiplied top and bottom by c so its easier to compute. then plugging it in: \[\lambda' = 0.144nm + (1240eV/(0.511*10^6 eV)) (1-\cos60)=0.145nm\] Since the electron is at rest to start, \[Pphoton=P'photon +Pe'\] That was teh first step I did

OpenStudy (anonymous):

Well, you're thinking about it wrong, it's not like conservation of energy. If you remember using Newton's laws, remember how we always break it into components for the force, we treat momentum the same way. So, lets look at the x direction, maybe you should draw a picture it will help you understand. \[\huge x: \frac{ hf }{ c } + 0 = \frac{ hf' }{ c }\cos(\phi)+p_{ex}\] \[\huge y: 0 + 0 = - \frac{ hf' }{ c } \sin (\theta) + p_{ey}\]

OpenStudy (anonymous):

Now you can solve for Pex and Pey and then use Pythagorus and tan ratio to find the angle.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

well using that I get \[4.6*10^-24 = 4.57*10^-24 \cos \phi+P'e\] \[P'ey=3.96*10^-24\]

OpenStudy (anonymous):

How do I move forward from there?

OpenStudy (anonymous):

I'm not checking your numbers, but it does look pretty good, so solve for Pex, and then we just have to use Pythagorean theorem, to find the resultant vector Pe.

OpenStudy (anonymous):

\[P_e = \sqrt{(P_{ex})^2+(P_{ey})^2}\]

OpenStudy (anonymous):

\[\sqrt{(4.6*10^-24 -4.57*10^-24 \cos \phi)^2+(3.96*10^-24)^2}\] What about the cos though?

OpenStudy (anonymous):

Which cos?

OpenStudy (anonymous):

Oh that's cos60

OpenStudy (anonymous):

cos60 and sin60 is what you're using since we just broke it into components

OpenStudy (anonymous):

For both I use an angle of 60? The electron should have a different angle than the photon?

OpenStudy (anonymous):

No, we're looking for the angle of the linear momentum of the electron just after the collision with the photon, that's the point of the question

OpenStudy (anonymous):

So we have to solve that still

OpenStudy (anonymous):

That's why we use \[\tan \theta = \frac{ P_{ey} }{ P_{ex} }\]

OpenStudy (anonymous):

Ah ok, well plugging all that in I get 4.59*10^-24

OpenStudy (anonymous):

Know the units? :)

OpenStudy (anonymous):

should be kgm/s

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

Ok, so thats our momentum for the electron just after the collision with the photon, now we just need the angle, as I showed above, solve for theta.

OpenStudy (anonymous):

=59.7 degrees! :D and its right! wow thanks soooo much for taking the time to help, I have been stuck on that problem for a long time!

OpenStudy (anonymous):

There we go :)

OpenStudy (anonymous):

Np ^.^, that's why I'm here right hehe